OpenStudy (anonymous):

how do you solve 3xsquared - 12x + 12 = 0

OpenStudy (anonymous):

Factor then find the zeros.

OpenStudy (anonymous):

can you show me?

OpenStudy (dumbcow):

Divide out the common factor of 3 first

OpenStudy (anonymous):

ok and then?

OpenStudy (anonymous):

E.g. 3 (x^2 - 4x + 4)

OpenStudy (dumbcow):

Notice its a perfect square...(x-2)^2 or look for factors of 4 which add up to -4

OpenStudy (anonymous):

\[x^2-2x-2x+4=0\] \[x(x-2)-2(x-2)=0\] \[(x-2)(x-2)=0\] x=2